for low speed ( laminar) flow through a circular pipe, the velocity distribution takes the form v=(beta/


1
Expert's answer
2021-05-25T10:11:48-0400

"v=B\dfrac{\Delta p}{\mu}(r_o^2-r^2)"



Dimension of "v=L\/T=LT^{-1}"

Dimension of "\mu=\dfrac{\tau}{\dfrac{du}{dy}}"

"\mu=ML^{-1}T^{-1}"

Dimension of "\Delta p=ML^{-1}T^{-2}"

Dimension of "r=L"

Substituting all values in first equation

"LT^{-1}=B\times\dfrac{ML^{-1}T^{-2}}{ML^{-1}T^{-1}}\times L^2"

"B=L^{-1}"

Therefore, dimension of B = L-1